OpenStudy (anonymous):

2x/16 +30=-1/x How do i do this? help please :/

hartnn (hartnn):

try multiplying 16x on both sides..that is to all terms....and simplify

OpenStudy (anonymous):

okay ill try that, thanks(:

OpenStudy (anonymous):

First you need to get all the Xs to one side: -30-15 = x-2x Then simplify: -45 = -x Since you have a negative on both sides, that will cancel out. Soooo: x = 45

OpenStudy (anonymous):

i got 8 and -8 for the awsner ????

OpenStudy (phi):

if you replace x with 8 in the equation, is the equation true, after simplifying?

OpenStudy (anonymous):

i tryed but then i got stuck

OpenStudy (anonymous):

x=-5,-3

OpenStudy (phi):

if the problem is \[ \frac{2x}{16}+30=\frac{-1}{x} \] I would notice that \[ \frac{\cancel{2}x}{\cancel{2}\cdot 8}+30=\frac{-1}{x} \] to simplify it a bit now multiply all the terms by 8x. what do you get? (write out the whole equation)

OpenStudy (phi):

to multiply, just write 8x* in front of every term on both sides of the equal sign

OpenStudy (anonymous):

is it suppose to be a 16x instead of just 16